OpenStudy (anonymous):

2. Fiona has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $25.10. If x represents the number of quarters in the jar, which equation can be used to find the number of quarters? (Points : 1) 0.25x + 0.05(x – 150) = 25.10 0.05x + 0.25(150 – x) = 25.10 0.05x + 0.25(x – 150) = 25.10 0.25x + 0.05(150 – x) = 25.10

OpenStudy (anonymous):

would it be A? that's what i think it is.

OpenStudy (jdoe0001):

if we make "x" the quarters, then we can say that \(\bf \text{x + "nickels amount" = 150 coins}\) and we can also say that \(\bf \text{"nickels amount" = 150-x}\) solving for "nickels amount" there are 25 cents in 1 quarter, and 5cents in 1 nickel so we can say that, the amount of quarters times 25 cents, "x", plus the "nickels amount" times 5 cents= $25.10 25.10 is really just 2,510 cents or pennies so so we can say that 25x + 5(x-150) = 2,510 "x" + (150-x) = 150

OpenStudy (jdoe0001):

25x + 5(x-150) = 2,510 #divide both sides by 100 to convert it from pennies to dollars

OpenStudy (anonymous):

thx:D

OpenStudy (jdoe0001):

yw